### v3

# The 1/3 Geometric Constant: Scale Invariance and the Origin of ’Missing Energy’ in 3D Quantum Fragmentation

Abstract
We report the discovery of a universal geometric constraint on the detection of kinetic energy release (KER) in three-dimensional quantum fragmentation. By investigating the dissociation of localized Slater-type orbitals, we demonstrate that the $4\pi r^{2}$ radial volume element inherently masks the majority of a system’s energy budget, imposing a fundamental peak-to-mean energy bound of $R_{E}<0.5$. We introduce a dimensionless scaling law, $M=\alpha\zeta/Q$, and prove that the resulting energy detection ratio is scale-invariant across twelve orders of magnitude, bridging the gap between atomic and subatomic physics. Remarkably, we identify a “Geometric Zero-Point” at $R_{E}\approx 0.33$, which precisely replicates the 7 eV “missing energy” anomaly observed in attosecond $H_{2}^{+}$ fragmentation benchmarks. Furthermore, we demonstrate that this $1/3$ ratio provides a robust geometric baseline for the historical average-to-endpoint discrepancy in beta decay. Our results suggest that a significant portion of what is historically categorized as “missing energy” may be a topological artifact of 3D quantum geometry rather than an exclusive signature of undetected particles. This work establishes a universal master curve for energy reconstruction and identifies a “detection crisis” in highly localized systems, where the true interaction energy becomes effectively invisible to peak-centric experimental calorimetry.
pacs:
32.80.-t,32.80.Rm,32.80.Fb
preprint: APS/123-QED
I Introduction
A central challenge in experimental physics is the reconstruction of the total energy budget from a detected momentum distribution. Across disparate scales—from the electron-volts of molecular dissociation to the mega-electron-volts of nuclear transitions—a persistent anomaly remains: the primary spectral peak often captures only a fraction of the available energy. Historically, this “missing energy” has been the catalyst for some of the most profound discoveries in physics, most notably the hypothesis of the neutrino.
The birth of the neutrino was necessitated by the observation that the average energy of electrons emitted in beta decay was approximately one-third of the total transition energy (the endpoint) Ellis and Wooster (1927). To preserve energy conservation, Pauli famously postulated an invisible particle to carry the balance.
Remarkably, a nearly identical $1/3$ discrepancy has emerged in modern attosecond calorimetry. In the Coulomb fragmentation of $H_{2}^{+}$, classical energy conservation and 1D models predict a kinetic energy release (KER) peak at $\approx 11$ eV Yue and Madsen (2013). However, high-precision experimental measurements and full-dimensional Time-Dependent Schrödinger Equation (TDSE) simulations consistently report a primary peak between 3344 eV Zhu (2020, 2021). This observed ratio of $R_{E}\approx 0.32$ almost perfectly mirrors the historical beta decay discrepancy, despite the absence of neutrinos in molecular fragmentation.
In this work, we demonstrate that these anomalies share a common, purely geometric origin. We prove that for any localized 3D ground-state wavefunction, the $4\pi r^{2}$ volume element inherently suppresses probability near the origin, sequestering the majority of the interaction energy in a high-momentum radial tail. We identify a universal 0.33 geometric landmark that emerges under a specific dimensionless scaling law, $M=\alpha\zeta/Q$, where $M$ is the particle mass, $\zeta$ the orbital localization, and $Q$ the repulsive charge.
Specifically, we show that as a system becomes more massive and more localized—as in the case of nucleons where $\zeta$ increases alongside MM—the detected peak energy remains a small, fixed fraction of the true energy. This allows us to establish a master curve for energy reconstruction, suggesting that much of the energy historically labeled as “missing” is actually a topological artifact of 3D quantum geometry. This framework provides a new tool for particle verification and discovery, shifting the focus from undetected particles to the intrinsic masking effects of the 3D volume element.
II Methods
III Theoretical Methodology
III.1 The Sudden Approximation and Potential Inversion
The core of our model rests on the sudden approximation, which is valid when a system’s Hamiltonian changes on a timescale $\tau_{perturb}$ much shorter than the characteristic response time of the quantum state:
$\tau_{perturb}\ll\frac{\hbar}{\Delta E}$ (1)
where $\Delta E$ represents the energy spacing of the system. Under this condition, the initial wavefunction $\psi_{0}(\mathbf{r})$ remains “frozen” at $t=0$, as there is insufficient time for the probability density to reorganize Makarov (2016); Eichler and Meyerhof (1995).
In the context of a Coulomb explosion, the potential undergoes an instantaneous transition from an attractive configuration to a repulsive one:
$V_{initial}(r)=-\frac{Z_{att}}{r}\longrightarrow V_{final}(r)=+\frac{Q}{r}$ (2)
This approximation effectively captures the dynamics of ultrafast processes such as multi-photon ionization by attosecond pulses, core-hole creation followed by Auger decay, or the sudden loss of strong-force confinement in nuclear physics.
III.2 Bohmian Local Energy Formulation
To determine the kinetic energy distribution of the resulting fragments, we utilize the de Broglie-Bohm formulation of quantum mechanics. The atomic units $\hbar=m_{e}=e=1$ are applied if not specfied. In this framework, the total energy $E(\mathbf{r})$ of a particle at a specific spatial coordinate $\mathbf{r}$ is defined by the sum of the repulsive potential and the local kinetic energy density $T_{local}(\mathbf{r})$:
$E(\mathbf{r})=T_{local}(\mathbf{r})+V_{final}(\mathbf{r})$ (3)
The local kinetic energy is derived from the Laplacian of the initial spatial wavefunction:
$T_{local}(\mathbf{r})=-\frac{1}{2M}\frac{\nabla^{2}\psi_{0}(\mathbf{r})}{\psi_{0}(\mathbf{r})}$ (4)
For the Slater-type orbitals (STOs) investigated in this work, defined by the radial form $R(r)=Nr^{n-1}e^{-\zeta r}$, the Laplacian in spherical coordinates yields the following analytical expression for the local kinetic energy:
$T_{local}(r)=\frac{1}{2M}\left(\zeta^{2}-\frac{2n\zeta}{r}+\frac{n(n-1)}{r^{2}}\right)$ (5)
Combining this with the repulsive potential, we obtain the conserved total energy for a trajectory starting at radius $r$:
$E(r)=\frac{1}{2M}\left(\zeta^{2}-\frac{2n\zeta}{r}+\frac{n(n-1)}{r^{2}}\right)+\frac{Q}{r}.$ (6)
III.3 Numerical Implementation and Statistical Weighting
The final observable energy distribution $P(E)$ is constructed by sampling the initial positions according to the 3D radial probability density:
$\rho(r)=4\pi r^{2}|\psi_{0}(r)|^{2}$ (7)
By mapping each sampled coordinate $r$ to its corresponding energy $E(r)$, we generate a histogram weighted by $\rho(r)$. Because each Bohmian particle conserves its total energy along its trajectory in a time-independent repulsive field, the energy distribution calculated at the instant of the potential flip ($t=0$) is identical to the asymptotic kinetic energy distribution measured by detectors at $t\to\infty$.
This methodology avoids the computational complexity of solving the Time-Dependent Schrödinger Equation (TDSE) while preserving the critical 3D geometric information—namely the $r^{2}$ volume weighting—that is lost in simplified 1D models.
IV Distribution property
To validate the predictive power of our 3D Bohmian-based model, we compared the resulting energy distributions $P(E)$ against full quantum simulations performed via the tRecX Scrinzi (2022); Zhu and Scrinzi (2020); Zhu (2021, 2020); Majety and Scrinzi (2015); Zielinski et al. (2016) software package. We utilized a standard 1s Slater orbital ($\zeta=1$,$l=0$) as the initial state and varied the repulsive Coulomb charge $Q$. We will atomic units
As shown in upper figure of Fig 1, the peak positions calculated by our method are in excellent agreement with the tRecX benchmarks across all charge values ($Q\in\{0,0.5,1\}$), the shape of our distribution is also similar to tRecX computation. The deviations in the distribution tails are attributed to the numerical constraints of grid-based solvers. Specifically, the truncation of the Coulomb potential and the use of polynomial quadrature grids in tRecX often struggle to resolve the singularity at $r\to 0$, whereas our analytical local-energy approach maintains precision in the near-nucleus region. This consistency in peak prediction confirms that our simplified method effectively captures the primary experimental signatures of the Coulomb explosion.
Figure 1: Comparison of normalized energy distributions P(E) between the present model (solid) and tRecX TDSE simulations (dashed). Results are shown for repulsion charges $Q=0,0.5$, and using an initial 1s hydrogenic orbital ($\zeta=1,l=0$). Peak positions show high consistency across all charge values. Parametric evolution of the energy distribution $P(E)$. (Middle) Variation with repulsive Coulomb charge $Q\in[0,2]$ for a fixed 1s orbital ($\zeta=1$,$l=0$). Both the peak position and spectral width increase with Q. (Lower) Variation with orbital exponent $\zeta\in[1,4]$ at a fixed charge $Q=1$. The distribution broadens more rapidly with increasing $\zeta$ compared to $Q$, highlighting the sensitivity of the KER to initial state localization.
To characterize the influence of the system’s physical parameters on the detected energy, we performed a systematic study varying the repulsive Coulomb charge $Q$ and the orbital exponent $\zeta$.
As illustrated in Fig. 1 (Upper), increasing the repulsive charge $Q$ leads to a systematic shift in the energy distribution $P(E)$ toward higher energies, where the peak of $P(E)$ is proportional to $Q$. This is accompanied by a broadening to some extend of the spectral width. Since the local energy is defined as $E(r)=T_{local}(r)+Q/r$, a larger $Q$ increases the potential gradient across the spatial extent of the orbital, thereby spreading the resulting energy values over a wider range. The shape of the curve could be fited by a radial distribution, as is illustrated in Sec. A in the appendix, suggesting inherently connection with the radial coordinates.
In Fig. 1 (Lower), we examine the effect of the orbital exponent $\zeta$ while maintaining a constant charge ($Q=1$). We observe that the peak position of $P(E)$ is also proportional to $\zeta$. As $\zeta$ increases, the wavefunction becomes more tightly bound, which simultaneously increases the local kinetic energy density $T_{local}\propto\zeta^{2}$ and places the particle density in regions of higher potential energy. Interestingly, the distribution broadens much more rapidly with respect to $\zeta$ than it does with respect to $Q$. This suggests that the internal ”quantum pressure” (local kinetic energy) of the initial state plays a more dominant role in determining the final energy dispersion than the external repulsive force itself.
Figure 2: Universal scaling of the peak-to-mean energy ratio $R_{E}=E_{\text{peak}}/\langle E\rangle$. (Upper) Variance with repulsive charge $Q$ at $\zeta=1$, showing monotonic convergence toward 0.5 as the repulsive potential dominates. (Middle) Variance with orbital exponent $\zeta$ at $Q=1$, showing a precipitous drop in detection efficiency for localized states. (Lower) Contour map of $R_{E}(\zeta,Q)$. The ratio is strictly bounded below 0.5, with the greatest discrepancy occurring for inner-shell analogs (high $\zeta$). High $\zeta$ values correspond to deeper initial potentials where the “quantum pressure” ($T_{\text{local}}\sim\zeta^{2}$) significantly skews the average energy relative to the experimental peak.
V Scaling Laws and the Energy Detection Discrepancy
A critical quantitative observation is found in the integrated energy values shown in the plot headers. In both cases, the peak energy $E_{\text{peak}}$ is significantly smaller than the total scattering energy $E_{\text{scatter}}$ (the integrated average $\langle E\rangle$). Specifically, for the $\zeta=1,Q=1$ case, the ”energy conservation” value $\langle E\rangle$ is driven higher by the potential energy density near the origin, whereas the experimental peak corresponds to the most probable radial position. This discrepancy validates our hypothesis that fragment detectors fundamentally probe the radial geometry of the initial state rather than the integrated energy mean.
To quantify the discrepancy between experimentally observed fragment signals and theoretical energy conservation, we define the energy detection ratio as:
$R_{E}=\frac{E_{peak}}{\langle E\rangle}$ (8)
where $E_{peak}$ is the mode of the distribution $P(E)$ and $\langle E\rangle$ is the integrated average (expectation value) of the total energy. This dimensionless ratio serves as a metric for the “skewness” of the distribution and the efficiency with which a detector captures the system’s total energy.
Figure 2 illustrates the systematic evolution of $R_{E}$ across a broad parametric space of repulsive charges ($Q$) and orbital exponents ($\zeta$).
V.1 Influence of Repulsive Charge $Q$
As shown in the upper panel of Figure 2, $R_{E}$ increases monotonically with the repulsive charge $Q$, asymptotically approaching a limit of $0.5$. In the high-$Q$ regime, the repulsive potential $V=Q/r$ dominates the local kinetic energy terms. This “classicalization” of the potential landscape shifts the peak of the distribution closer to the statistical mean. However, the fact that $R_{E}$ remains strictly bounded below $0.5$ underscores that the 3D quantum spatial spread of the Slater orbital is an inescapable constraint on the fragment energy spectrum.
V.2 Orbital Exponent $\zeta$: Inner-Shell vs. Outer-Shell Dynamics
The middle panel of Figure 2 demonstrates a precipitous decline in $R_{E}$ as the orbital exponent $\zeta$ increases. In the framework of Slater-type orbitals, $\zeta$ serves as a proxy for the effective nuclear charge and the principal quantum number ($\zeta\approx Z_{eff}/n$). Consequently, small $\zeta$ values represent diffuse outer-shell (valence) orbitals, while large $\zeta$ values characterize highly localized inner-shell (core) orbitals.
As $\zeta$ increases, the wavefunction contracts, and the local kinetic energy density scales as $T_{local}\propto\zeta^{2}$. This quadratic scaling of the “quantum pressure” near the nucleus drives the integrated average energy $\langle E\rangle$ to values much higher than the peak energy $E_{peak}$, which is primarily determined by the most probable radial position $r_{mp}=1/\zeta$.
Our findings suggest that the reliability of fragment kinetic energy as a proxy for interaction strength is shell-dependent. For outer-shell Coulomb explosions (low $\zeta$), $R_{E}$ is relatively high, indicating that the detected peak provides a reasonably transparent window into the system’s potential energy. Conversely, inner-shell events (high $\zeta$) are subject to a massive quantum geometric shift; for $\zeta\approx 5$, the detector may observe as little as 15% of the integrated energy mean. Furthermore, as $\zeta$ scales with the particle mass, the extreme localization of nucleons in nuclear physics implies a significantly suppressed $R_{E}$ ratio, suggesting that the detected kinetic energy in nuclear fragmentation may represent only a small fraction of the total potential energy stored in the initial 3D quantum state.
V.3 Universal Contour Map and Numerical Stability
The global contour plot in Figure 2 (lower panel) synthesizes these trends, confirming that the $R_{E}<0.5$ boundary is maintained across the entire $(\zeta,Q)$ parameter space. Minor fluctuations observed at the parameter extremes are attributed to numerical truncation errors at the $E\to 0$ and $E\to\infty$ tails of the distribution. Despite these limits, the global trend is unambiguous: the deeper a particle is “buried” in the atomic potential (larger $\zeta$), the more the 3D quantum geometry masks the true interaction energy from classical detection.
V.4 The Dimensionless Master Curve and Scaling Symmetry
The emergence of the 0.33 landmark is rooted in a fundamental scaling symmetry of the three-dimensional fragmentation dynamics. To characterize this, we examine the peak-to-mean ratio $R_{E}$ across the parametric landscape of mass ($M$), orbital localization ($\zeta$), and repulsive charge ($Q$). As demonstrated in Fig. 3, the ratio is governed by a single dimensionless coupling constant $\alpha$ when satisfying:
$\alpha=\frac{MQ}{\zeta}.$ (9)
Figure 3: Universal Scaling and the 0.33 Geometric Landmark. (Top) $R_{E}$ vs. $Q$ for the case $\zeta=Q,M=1,n=1$, showing a stable plateau at 0.33 (dashed line). (Middle) $R_{E}$ vs. $\zeta$ for the case $Q=1,M=\zeta,n=1$, exhibiting identical stabilization at 0.33. (Bottom) 3D contourf representation of the parameter space $(\zeta,Q)\in[1,5]$ with the mass scaling $M=\alpha\zeta/Q$. The emergence of three parallel, invariant planes for $\alpha=0.5$ ($R_{E}\approx 0.26$), $\alpha=1.0$ ($R_{E}\approx 0.33$), and $\alpha=2.0$ ($R_{E}\approx 0.39$) confirms that the energy detection efficiency is a pure function of the dimensionless scaling factor $\alpha$.
By substituting our proposed scaling relation (Eq. 9) into the kinetic term, we obtain:
$\langle T\rangle\propto\frac{\zeta^{2}}{\alpha\zeta/Q}=\frac{1}{\alpha}\zeta Q.$ (10)
Crucially, under this transformation, both the kinetic and potential energy components scale linearly with the product $\zeta Q$. This synchronization ensures that the relative weights of the energy components—and thus the geometric morphology of the resulting distribution P(E)P(E)—depend solely on the dimensionless coupling constant $\alpha$. This is also illustrated in the upper figure of Fig. 3, where we plot the $P(E)$ curve with scaled parameter $\frac{E}{Q\zeta}$. We found the base curve with variant $Q$ where $\zeta=M=n=1$, matches eactly with curves of variant $\alpha$ (the parameters satisfying Eq. 9). As can be illustated with the exact match of base curve $Q=1$ with $\alpha=1$ ($Q=2,\zeta=1$ and $Q=1,\zeta=2$), and base curve $Q=2$ with $\alpha=2$ ($Q=2,\zeta=1$ and $Q=1,\zeta=2$). The scaled shape of $P(E)$ is controlled by the parameter $\alpha$, where higher $\alpha$ values result in a narrower distribution with a constant peak position. This behavior provides an intuitive representation of the $\alpha$-dependent scaling ratio. By using this distribution as a baseline to fit experimental data, one can uncover fundamental properties like $M$, $\zeta$ and $Q$. We demonstrate this utility later by comparing our model to the beta decay spectrum.
One step further, this shows $R_{E}$ remains invariant along the $M\propto\zeta$ trajectories: the system undergoes a self-similar expansion in energy space where the $4\pi r^{2}$ masking effect is preserved as a constant geometric ratio. The invariance observed in the parallel planes of Fig. 3 (lower figure) can be derived by examining the homogeneous scaling properties of the system’s energy components. For a 1s Slater orbital, the expectation value of the potential energy scales linearly with the orbital exponent, $\langle V\rangle\propto Q\zeta$, while the expectation value of the kinetic energy (quantum pressure) scales quadratically with the exponent and inversely with the mass, $\langle T\rangle\propto\zeta^{2}/M$.
To determine the universal values of $R_{E}$ across the scaling parameter $\alpha$, we utilize the invariance property to map the ratio onto our previously computed $Q$-dependence results (where $M=\zeta=n=1$) from Fig. 2:
$R_{E}(\alpha)=R_{E}(Q)\rvert_{M=\zeta=1}.$ (11)
The figure is also re-ploted in Fig. 4,
Figure 4: $R_{E}$ vs. $\alpha$ for the case, same as computations with baseline for Q $M=\zeta=n=1$.
with numerical behavior of $R_{E}$ is thus characterized by the following limits:
$\lim_{\alpha\to 0}R_{E}=0,\quad\lim_{\alpha\to\infty}R_{E}=0.5.$ (12)
When $\alpha=1$, the system achieves a “geometric equilibrium” where the $4\pi r^{2}$ volume element sequesters exactly two-thirds of the total energy in the high-momentum radial tail, yielding the 0.33 landmark. As $\alpha$ decreases (physically representing high $Q$ or low $M$), the ratio monotonically climbs toward the 0.5 universal bound, a regime where the repulsive field dominates the initial-state quantum pressure.
Conversely, as $\alpha$ increases (high localization $\zeta$), $R_{E}$ decays precipitously, falling below 0.2 as $\zeta\to 5$ and asymptotically approaching 0. This signals a profound “detection crisis” where the vast majority of the interaction energy becomes experimentally invisible to peak-centric calorimetry. This confirms that for sufficiently localized states—such as core-shell electrons or compressed nucleons—the true interaction energy is effectively masked by the topological constraints of 3D space. Consequently, the mapping $R_{E}(\alpha)$ serves as a universal master curve for energy reconstruction across all scales of 3D quantum fragmentation.
VI Principal Quantum Number Effects and Multi-Peak Phenomena
Figure 5: Influence of the principal quantum number $n$ on the energy distribution $P(E)$ for $Q=1,\zeta=1$. While the primary peak position remains consistent with the $n=1$ state, a secondary high-energy peak emerges for $n=2$ and $n=3$. The differences in the high-energy cutoff for varying $n$ are attributed to limited grid spacing in the numerical treatment. The growth of this secondary feature explains the extra peaks observed in experimental KER spectra where higher shells are populated.
Beyond the ground state ($n=1$), higher principal quantum numbers ($n>1$) introduce additional radial nodes and qualitative changes to the topology of the energy distribution $P(E)$. Figure 5 illustrates the computed spectra for $n=1,2,$ and $3$, maintaining fixed parameters of $Q=1$ and $\zeta=1$ with $M=1$.
Our simulations reveal that while the position of the primary (low-energy) peak remains relatively stable as we move to higher shells, the increase in the principal quantum number $n$ triggers the emergence of a secondary peak at significantly higher energies. The magnitude of this secondary feature scales positively with $n$.
For excited states ($n>1$), the local kinetic energy $T_{local}(r)$ incorporates terms related to the curvature of the additional radial nodes. Physically, this redistributes probability density into regions of higher local kinetic energy, creating a multi-modal distribution. In the high-$n$ limit, the secondary peak can undergo a dominance shift, eventually surpassing the primary peak to become the most prominent feature of the distribution.
We note that the varying lengths of the energy distribution curves for $n=1,2,$ and $3$ in our plots are an artifact of the numerical treatment. Specifically, the high-energy truncation points are sensitive to the grid spacing used during discretization, which limits the resolution of the extremely high-energy tails as the spatial complexity of the orbital increases.
VII Evidence Across Atomic and Subatomic Scales
VII.1 Empirical Validation: The $H_{2}^{+}$ Energy Discrepancy
Figure 6: Proton KER spectra for $H_{2}^{+}$ dissociative ionization. Dashed and solid lines represent 1D (taken from Ref. Yue and Madsen (2013)) and 3D computations, respectively. Triangle markers denote experimental data taken from Ref. Pavičić et al. (2005). The black curves represent ground-state ($n=1$) contributions, while red curves include contributions from populated excited states.
The physical validity of the 0.33 geometric landmark is most strikingly evidenced by the long-standing energy discrepancy in attosecond $H_{2}^{+}$ fragmentation. As shown in Fig. 6, traditional 1D semiclassical models Yue and Madsen (2013) predict a KER peak between 10101313 eV, significantly overestimating the experimental observation. In contrast, our 3D simulations (solid black line), performed via the tRecX quantum framework, yield a primary peak near $3$ eV, in excellent agreement with experimental calorimetry.
Our model replicates this $\approx 7$ eV shift without requiring external field effects or complex multi-photon dynamics. Although the $H_{2}^{+}$ system lacks perfect central symmetry and involves fragments with significant mass ($M_{p}\approx 1836$ a.u.), the geometric constraint remains the dominant physical driver. Because the initial $1\sigma_{g}$ state is highly localized, the $4\pi r^{2}$ volume weighting effectively sequesters the majority of the potential energy in the high-momentum radial tail. Consequently, the most probable detected energy ($E_{\text{peak}}$) represents only one-third of the integrated mean energy ($\langle E\rangle\approx 11$ eV), confirming that the “missing energy” in molecular physics is a direct manifestation of 3D quantum geometry.
Furthermore, when excited states ($n>1$) are involved, our 3D model identifies a secondary high-energy feature between 4455 eV (red curves in Fig. 6). This secondary peak, which is entirely absent in 1D approximations, serves as a signature of the radial nodal structures in the initial state. While excited-state populations are typically secondary to the $n=1$ signal, their presence explains the characteristic broadening and high-energy structured tails observed in experimental KER spectra.
VII.2 The Geometric Origin of Missing Energy in Subatomic Physics
Figure 7: Comparison of the historical beta decay spectrum of Radium E (210Bi), adapted from Ref. Ellis and Wooster (1927) (triangle markers), and our theoretical $P(E)$ distribution with $\alpha=0.16$ (red solid line). The spectrum exhibits the characteristic 1/3 ratio between mean and endpoint energy, consistent with our predicted geometric landmark.
The identification of the $M=\alpha\zeta/Q$ scaling law provides a provocative new lens for interpreting energy anomalies in nuclear physics. Historically, the continuous electron spectrum observed in beta decay—characterized by an average energy $\langle E\rangle\approx 1/3E_{\text{max}}$—was the primary catalyst for the neutrino hypothesis Ellis and Wooster (1927). Notably, the morphology of the beta spectrum (Fig. 7) is remarkably well-replicated by our model in the low-to-mid energy range ($0.5E_{\text{max}}$) using a coupling constant of $\alpha=0.16$. This value corresponds to a high-localization ($\zeta$) regime, consistent with the high atomic number ($Z$) of Radium.
The divergence of the fit at higher energies is expected, as this region enters the relativistic regime where the non-relativistic Laplacian used in our current $P(E)$ derivation reaches its limit of applicability. In this regime, the operator overestimates the energy—driven, causing the fit to curve above the Beta Decay baseline, as illustrated in the figure. Furthermore, the nuclear $Q$-value imposes a hard phase-space cutoff at the endpoint energy that truncates the theoretical high-energy tail. Despite these high-energy corrections, our results suggest that the $R_{E}\approx 0.33$ ratio represents a fundamental geometric baseline for energy detection in localized 3D systems. This geometric constraint potentially offers a unified explanation for perceived missing energy across diverse regimes, from Auger-induced Coulomb explosions to the quenching factors observed in dark matter direct detection Levin and others (1991); Aprile and others (XENON Collaboration) (2023).
VII.3 Unified Mechanism of Spectral Broadening and Orbital Contraction
The spectral distribution $P(E)$ is critically governed by the localization parameter $\zeta$. As illustrated in our parametric study (Fig. 1), an increase in $\zeta$ induces a rapid broadening of the high-energy tail. In high-$Z$ nuclei, the proximity of inner-shell electrons effectively scales the $\zeta$ parameter, leading to the spectral broadening observed in traditional nuclear studies Wilkinson (1990).
For superheavy nuclei ($Z>100$), the extreme localization (high $\zeta$) increases the probability of detecting ultra-high energy fragments. This phenomenon aligns with experimental observations of relativistic orbital contraction Henning (2012). While currently interpreted via Fermi theory Pershina (2016), our model suggests these relativistic shifts are consistent with the geometric pressure inherent to high-$\zeta$ localized states, providing a scale-invariant link between atomic and relativistic nuclear dynamics.
VIII Conclusion
In this work, we have demonstrated that the persistent discrepancy between theoretical energy budgets and experimental peaks in quantum fragmentation is not merely a technical limitation, but a fundamental topological artifact of three-dimensional wavefunctions. By utilizing localized Slater-type orbitals and full-dimensional TDSE benchmarks, we have established that the $4\pi r^{2}$ volume element inherently masks a significant portion of a system’s kinetic energy, imposing a universal peak-to-mean energy bound of $R_{E}<0.5$.
The discovery of the scale-invariant law, $M=\alpha\zeta/Q$, provides a unified framework for energy reconstruction across twelve orders of magnitude. We have shown that the observed 1/3 ratio in $H_{2}^{+}$ dissociative ionization is a direct manifestation of this geometric constraint, replicated here without the need for complex multi-photon dynamics or external field corrections. Most significantly, the alignment between this “Geometric Zero-Point” and the historical average-to-endpoint discrepancy in beta decay suggests that 0.33 is a universal landmark of 3D detection.
Our findings offer a cautionary note for the interpretation of “missing energy” in particle and nuclear physics. While the neutrino and other non-interacting particles are essential components of modern theory, our results reveal that the baseline for energy detection in localized subatomic states is inherently skewed by the topology of space itself. Because the $1/3$ ratio persists—and can even decay toward zero—as particles become more massive and localized, this geometric masking must be decoupled from signatures of invisible flux. This perspective shifts the focus of future calorimetry from the search for undetected mass to the correction of the intrinsic geometric filters of the three-dimensional quantum world.
Appendix A Fiting curve of $P(E)$
Figure 8: Analytical fitting and energy comparison of P(E). (Upper) Results for $\zeta=1,Q=1,l=0$; (Lower) Results for $\zeta=5.2,Q=1.6,l=0$. The r-fit (green) and Gamma fit (cyan) show the lowest error (¡2%), identifying the radial nature of the distribution. $E_{p}$ and $E_{k}$ denote integrated potential and kinetic energies, while Escatter represents the total energy in the repulsive regime. In all cases, $E_{peak}$ remains substantially lower than Escatter.
To identify a universal analytical form for the fragment energy spectrum, we performed a regression analysis on the energy distribution P(E) using several statistical models. Fig. 8 illustrates the results for a ground-state hydrogenic orbital ($\zeta=1,Q=1$) and a more localized, higher-charge configuration ($\zeta=5.2,Q=1.6$).
We compared the computed data (blue) against three candidate distributions: a radial-type distribution (r-fit), a Gamma distribution, and a $\chi^{2}$ distribution. As shown in Fig. 2, the r-fit (green) and the Gamma distribution (cyan) produce the highest fidelity, with a residual error of less than 2%. The near-perfect overlap between these two functions confirms that the energy distribution essentially obeys a radial distribution law.